# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v1

@@ Abstract
In the quest to facilitate the deep intelligence of Large Language Models (LLMs) accessible in final-end user-bot interactions, the art of prompt crafting emerges as a critical yet complex task for the average user. Contrast to previous model-oriented yet instruction-agnostic Automatic Prompt Optimization methodologies, yielding polished results for predefined target models while suffering rapid degradation with out-of-box models, we present Free-form Instruction-oriented Prompt Optimization (FIPO). This approach is supported by our large-scale prompt preference dataset and employs a modular fine-tuning schema. The FIPO schema reimagines the optimization process into manageable modules, anchored by a meta prompt that dynamically adapts content. This allows for the flexible integration of the raw task instruction, the optional instruction response, and the optional ground truth to produce finely optimized task prompts. The FIPO preference dataset is meticulously constructed using the optimal and suboptimal LLMs, undergoing rigorous cross-verification by human experts and analytical models. Applying the insights from the data with Tulu2 models and fine-tuning strategies, we validate the efficacy of FIPO schema across five public benchmarks [^1] Codes, data and scripts are here: https://github.com/LuJunru/FIPO_Project..
@@
[…]
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated their exceptional intelligence Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) by excelling in various standard benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). Nevertheless, LLMs can not fully showcase their potential without carefully crafted prompts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b), especially when engaging with non-experts in final-end user-bot interactions.
Figure 1: An example of comparison between other LLM-based discrete APO strategies with our FIPO.
In recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, fpcusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce gradient-free, straightforward strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)). These approaches typically involve iterative processes between a prompt generator LLM and an evaluator LLM. Inspired by GAN Goodfellow et al. (2020), the generator iteratively explores and tests superior prompts, while the evaluator provides feedback on the optimized prompts and the testing results. However, despite their effectiveness, several drawbacks exist: (1) There are unavoidable instability in API-reliant frameworks; (2) The ad-hoc optimization tends to be model-centric, leading to performance degradation when tested with out-of-box models; (3) Many LLM-based discrete APO methods focus on the optimization of instruction-agnostic task descriptions, whereas optimizing instruction-oriented sections is more critical for assisting smaller LLMs.
To address the challenges and break free from the constraints imposed by pre-defined optimization templates in some existing methods Zhou et al. (2023), we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). Figure 1 illustrates a comparison between two mainstream LLM-based discrete APO strategies and our FIPO. On the left-hand is an example of In-Context-Learning Dong et al. (2023) prompt, consisting of a leading introduction sentence, two separate examples, and the testing task instruction “Calculate the average value of the list”. In particular, the two in-context examples follow a consistent format: a one-line question, a line-by-line Chain-of-Thought Wei et al. (2022) solution, and a final answer reporting template (“The answer is X”). It is worth mentioning that such few-shot template, is widely-accepted as the official testing format in popular benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Hendrycks et al. (2021).
In the right-hand section of Figure 1, we present a list of optimized prompts generated through various methods, arranged from top to bottom. The top-right method focuses on optimizing the top leading sentence. We define this category as “Instruction-agnostic” Optimization. In contrast, the middle-right approach, labeled as “Template-based” Optimization, involves incorporating specific guiding sentences into the entire prompt, accopmanied with a series of blank tokens that are filled in to optimize the testing instruction. Our proposed “FIPO” method, listed in the right-bottom part, avoids using preliminary templates and only considers the testing instruction for “Instruction-oriented” Optimization. Among all the optimized prompts, FIPO stands out as the only method capable of enhancing the testing instruction with embedded chain-of-thought guidance, instructing downstream models to “find the sum” and “divide it by the total number of elements” to more effectively complete the “Calculate the average value” task.
Our contributions can be summarized as follows:
  • We introduce our research of free-form instruction-oriented prompt optimization (FIPO), and subsequently devise a modular schema for in-depth investigation.
@@
  • Based on the schema, we compile a large-scale prompt preference dataset, and explore with several mainstream fine-tuning strategies: Supervised Fine-tuning, Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). We further develop an Iterative Preference Learning (IPL) pipeline for self-rewarding preference optimization.
@@
  • We evaluate the efficacy and adaptability of fine-tuned FIPO optimizer across various downstream benchmarks, and obtain significant improvements compared with previous LLM-powered discrete APO methods.
2 Related Work
Two related topics are involved: Automatic Prompt Optimization and Preference Optimization.
Figure 2: Harnessing data from a suboptimal LLM GPT-3.5 and an optimal LLM GPT-4.
2.1 Automatic Prompt Optimization
2.2 Preference Optimization for LLMs
Preference Optimization has been regarded as the key to the final mile of Artificial General Intelligence (AGI) Pan et al. (2023b). Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). A series of following work of DPO are further presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024).
@@ 3 Methodology
In this section, we start with task formulation, then introduce our schema, data and training strategies.
@@ 3.1 Task Formulation
Named as Free-form Instruction-oriented Prompt Optimization (FIPO), our approach follows the end-to-end text generation process. A neural optimizer model $\bm{M_{o}}$ is tasked with generating an optimized prompt $\bm{\hat{x}^{o}}=\operatorname*{argmax}_{\bm{M_{o}}}p(\bm{\hat{x}^{o}}|\bm{x}^{r},\bm{\hat{y}}^{r},\bm{y}^{r})$, based on the given raw task instruction $\bm{x}^{r}$, raw instruction response $\bm{\hat{y}}^{r}$, and ground truth $\bm{y}^{r}$. The raw instruction response $\bm{\hat{y}}^{r}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{r}|\bm{x}^{r})$ is generated from the raw task instruction $\bm{x}^{r}$ using a neural generator model $\bm{M_{g}}$. The ultimate target is to obtain a superior instruction response $\bm{\hat{y}}^{o}\succeq\bm{\hat{y}}^{r}$, where $\bm{\hat{y}}^{o}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{o}|\bm{\hat{x}}^{o})$. The raw instruction response $\bm{\hat{y}}^{r}$ and ground truth $\bm{y}^{r}$ are only provided during the training phase.
Figure 3: The overall visualization of our fine-tuning strategies.
3.2 Modular Schema and Preference Dataset
In this context, our objective is to develop a modular schema that ensures flexibility in content management, addressing potential exposure bias between the training and the inference phases. Simultaneously, we decide to distill the refined prompt optimization capabilities from prominent yet proprietary LLMs, instead of directly integrating them in an ad-hoc way. Figure 2 illustrates our research schema and how we simply harness contrastive optimized prompts with GPT-3.5 and GPT-4 APIs. We first establish a modular schema as shown in the middle. The meta optimizer prompt takes the raw task instruction $\bm{x}^{r}$ as mandatory input, while considering raw instruction response $\bm{\hat{y}}^{r}$ and ground truth response $\bm{y}^{r}$ as optional. Additional descriptive meta information is then appended for clarity.
@@ 3.2 Modular Schema and Preference Dataset
The raw task instruction $\bm{x}^{r}$ at the top is sampled from the Alpaca dataset Taori et al. (2023), which contains 52K diverse instructions and corresponding responses $\bm{\hat{y}}^{r}$ generated by the Text-davinci-003 API. We also collect another GPT-4-generated response for the Alpaca dataset from publicly available literature Ghosal et al. (2023). Although there are no official ground truth responses (e.g., from human experts) for the Alpaca data, we consider GPT-4 responses as the ground truth response $\bm{y}^{r}$, given its demonstrated analytical capabilities comparable to humans Pan et al. (2023a). Afterwards, we populate our meta-optimizer prompt with the aforementioned collected instruction $\bm{x}^{r}$, as well as two responses $\bm{\hat{y}}^{r}$ and $\bm{y}^{r}$, and then respectively apply GPT-3.5 and GPT-4 APIs to generate contrastive optimized prompt labels $\bm{x}^{o-}$ and $\bm{x}^{o+}$ in a zero-shot manner. As shown at the bottom of Figure 2, GPT-4 offers a more pedagogical step-by-step optimization compared to GPT-3.5.
@@ 3.3 Fine-tuning Strategies
We introduce our fine-tuning strategies in Figure 3, consisting of an initial step of dataset diversification followed by strategic fine-tuning.
Dataset Diversification. As illustrated in the left hand of Figure 3, in accordance to our modular schema, we evenly split the original 30k samples as eight types depending on the existence of raw instruction response $\bm{\hat{y}}^{r}$ and ground truth response $\bm{y}^{r}$, as well as a format condition “generation” or “multi-choice”. Robustly fine-tuning neural optimizer models have to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure divergence between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The original responses $\bm{\hat{y}}^{r}$ and $\bm{y}^{r}$ are modified as candidates adhere to the raw task instruction $\bm{x}^{r}$. We therefore set “A” and “B” as new task response and ground truth response [^2] More details of data diversification are in Appendix B..
@@
On the other hand, the right middle shows a very recent yet popular and effective methodology: Preference Optimization, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). Preference Optimization takes contrastive training labels $\bm{x}^{o-}$ and $\bm{x}^{o+}$ as supervision. One of the core differences from Preference Optimization and SFT is the former one not only encourage the generation of optimal preference, but also dampen the generation of suboptimal preference:
$L_{DPO}(M_{o})=-\mathbb{E}_{(x^{r},\hat{y}^{r},y^{r},x^{o+},x^{o-})\sim D}[log\sigma(\beta\Delta)]$ (2)
$L_{IPO}(M_{o})=-\mathbb{E}_{(x^{r},\hat{y}^{r},y^{r},x^{o+},x^{o-})\sim D}[\Delta-\frac{1}{2\beta}]^{2}$ (3)
$\Delta=log\frac{M_{o}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}-log\frac{M_{o}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}$ (4)
where $\bm{\beta}$ is a hyperparameter target. $\bm{M_{ref}}$ refers to the reference model, which is a frozen copy of initial $\bm{M_{o}}$. The equations indicate that IPO is a regularized version of DPO as it limits the optimization range with squares.
Additionally, inspired by recent super-human Reinforcement Learning from AI Feedback (RLAIF) Lee et al. (2023); Anthropic (2022); Yuan et al. (2024), we develop a Iterative Preference Learning (IPL) strategy for self-rewarding preference optimization. After each iteration of preference optimization, we ask the optimizer itself to determine if it successfully generate a superior prompt $\bm{x}^{r+}$ and response $\bm{\hat{y}}^{r+}$, and if so, to automatically replace the previous inferior prompt $\bm{x}^{r}$ and response $\bm{\hat{y}}^{r}$, leading to more rigorous training:
$L_{IPL}(M_{o})=-\mathbb{E}_{(x^{r+},\hat{y}^{r+},y^{r},x^{o+},x^{r})\sim D}G(\Delta)$ (5)
$x^{r+}=M_{o}(x^{r})$ (6)
$x^{r+}=\left\{\begin{array}[]{l}x^{r+},\quad M_{j}(x^{r+},y^{r})\succeq M_{j}(x^{r},y^{r})\\ x^{r},\quad otherwise\end{array}\right.$ (7)
$\hat{y}^{r+}=M_{g}(x^{r+})$ (8)
where $G(*)$ denotes as either IPO or DPO loss style. $\bm{M_{o}}$, $\bm{M_{j}}$ and $\bm{M_{g}}$ refers to the prompt optimizer, the prompt discriminator and the response generator, respectively. However, $\bm{M_{o}}\equiv\bm{M_{j}}\equiv\bm{M_{g}}$ in our settings [^3] Algorithmic introduction of IPL is in Appendix C.. We increment instruction following and discrimination data to maintain its judgement and answering abilities, shown in Appendix B.
The ultimate target of APO lies in the performance enhancement with downstream generators $\bm{M_{g}}$, using final optimized prompt $\bm{\hat{x}^{o}}$ to obtain a superior response $\bm{\hat{y}}^{o}\succeq\bm{\hat{y}}^{r}$. Herein, we evaluate our fine-tuning strategies across a variety of public benchmarks with various downstream generators.

### v2

@@ Abstract
When the quality of naive prompts is carefully optimized by human experts, the task performance of large language models (LLMs) can be significantly improved. However, expert-based prompt optimizations are expensive. Herein, some works have proposed Automatic Prompt Optimization (APO), to optimize naive prompts according to task outputs of given in-box testing models, with the help of advanced LLMs (e.g., GPT-4) in an ad-hoc way. Although effective, existing schemes suffer from poor generalization ability and privacy risk. To this end, we collect the first large-scale Prompt Optimization Preference dataset (POP), fine-tune offline local LLM-based optimizers, then fairly test with various downstream models. Our method allows accurate optimization of the core task instruction part within the naive prompt in a model-agnostic manner, and thus is named Free-from Instruction-oriented Prompt Optimization (FIPO). In specific, FIPO uses a modular APO template that dynamically integrate the naive task instruction, optional instruction responses, and optional ground truth to produce finely optimized prompts. The POP dataset is meticulously constructed using advanced LLMs, undergoing rigorous cross-validation by human experts and analytical models. Leveraging insights from the data with Tulu2 models and diverse fine-tuning strategies, we validate the efficacy of FIPO framework across five public benchmarks and three testing models [^1] Check codes and data here: https://github.com/LuJunru/FIPO_Project..
@@
Junru Lu1, Siyu An2, Min Zhang3, Yulan He4, Di Yin2, Xing Sun2 1University of Warwick, 2Tecent YouTu Lab, 3Westlake University, 4King’s College London 1junru.lu@warwick.ac.uk, 3zhangmin@westlake.edu.cn, 4yulan.he@kcl.ac.uk 2{siyuan, endymecyyin, winfredsun}@tencent.com
@@ 1 Introduction
Large Language Models (LLMs) show their intelligence Zhao et al. (2023a); Yang et al. (2023c); Achiam et al. (2023) on various benchmarks Cobbe et al. (2021); Suzgun et al. (2023); Bisk et al. (2020); Huang et al. (2019); Hendrycks et al. (2021). However, the task performance of LLMs depends on the given task prompt. LLMs may not be able to give the correct answer based on the naive task prompt, but after careful optimization by human experts, it can perform the same task well under the guidance of high-quality prompts Wei et al. (2022); Kojima et al. (2022); Yang et al. (2023b).
Figure 1: Online ad-hoc APO vs. our Local End2end FIPO. Although both leverage advanced LLMs (e.g., GPT-4), in FIPO, we design a local training pipeline and completely removed the reliance on any in-box model generators.
Obviously, expert-based prompt optimizations are expensive. Therefore, in recent years, Automatic Prompt Optimization (APO) has emerged as a prominent area of research. Discrete APO is one of the popular strategies, focusing on identifying optimal combinations of discrete tokens to serve as optimized prompts van de Kar et al. (2022); Yuan et al. (2021); Jiang et al. (2020); Pryzant et al. (2023). Particularly, there has been significant interest in LLM-based discrete APO Zhou et al. (2023); Do et al. (2023); Wang et al. (2023a), which introduce ad-hoc strategies leveraging leading API-accessible LLMs (e.g., GPT-4 Achiam et al. (2023)).
These APO approaches typically involve iterative optimization between an in-box testing generator $\bm{M}_{g-in}$ and an advanced optimizer $\bm{M}_{o-api}$. As illustrated in the upper half of Figure 1, the generator $\bm{M}_{g-in}$ first provide answers upon the naive prompt $\bm{x}^{n}$: “Calculate the average value of the list”, and then the optimizer $\bm{M}_{o-api}$ provides rational feedback and several upgraded prompt candidates $\{\bm{x}^{o}\}$. This above process can be repeated several times. Afterwards, a high-quality optimized prompt $\bm{x}^{o}$ is finally tailored for the in-box generator $\bm{M}_{g-in}$, which can be ensured to produce the desired response: “The answer value is 44.25”. However, despite effective, several drawbacks exist: (1) Privacy risk. The entire online optimization is exposed to external LLM services, which poses a privacy risk; (2) Poor Generalization Ability. The ad-hoc optimization is model-centric, as it runs with instant testing responses from the in-box generator $\bm{M}_{g-in}$, leading to performance degradation when tested with out-of-box generators $\bm{M}_{g-out}$. For instance, a wrong answer is inferred by the out-of-box generator $\bm{M}_{g-out}$: “44.1”.
To address the existing challenges, we introduce Free-form Instruction-oriented Prompt Optimization (FIPO). The bottom half of Figure 1 illustrates our FIPO. Unlike the online ad-hoc APO approach, FIPO directly fine-tunes the general local optimizer $\bm{M}_{local}$, and apply over any testing generator $\bm{M}_{g}$. In specific, we first design a meta template towards universal APO (Figure 2), and thereby collect 30,000 prompt optimization preference examples from advanced optimizer $\bm{M}_{o-api}$ (§ 2.3). We demonstrate the reliability of this data through multiple cross-validation methods (Table 1). On this basis, we then apply with mainstream end-to-end fine-tuning strategies for obtaining a valid local optimizer $\bm{M}_{o-local}$ (Figure 3). Particularly, in the Figure 1’s example, our local optimizer $\bm{M}_{local}$ provides insightful hints in the optimized prompt $\bm{x}^{o}$: “calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list”. Such stepwise guidance generally enhance the answer quality of any testing generator $\bm{M}_{g}$. Our contributions are as follows:
  • (1)
    We denote the drawbacks of previous online ad-hoc APO, and then introduce our local free-form instruction-oriented prompt optimization (FIPO). We design a generic APO template that allows the optimizer to unpack the pre-built testing generators, which is a key component for former methods.
  • (2)
@@
    Based on the template, we compile a 30k large-scale Prompt Optimization Preference dataset (POP), and explore with several mainstream fine-tuning strategies: Supervised Fine-tuning (SFT), Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). We further develop an Iterative Preference Learning (IPL) pipeline for self-enhanced preference optimization.
  • (3)
@@
    We evaluate the efficacy and adaptability of fine-tuned FIPO optimizer across five downstream benchmarks and three diverse generators, and obtain superior results compared to previous online ad-hoc APO methods.
Figure 2: Step 1 and 2 of FIPO: (1) Design a meta template towards universal APO; (2) Collect 30,000 large-scale prompt optimization preference exemplars from a suboptimal LLM GPT-3.5-turbo and an optimal LLM GPT-4.
@@ 2 Methodology
In this section, we start with task formulation (§ 2.1), then introduce our meta template towards universal APO (§ 2.2), collected POP data (§ 2.3) and training strategies (§ 2.4).
@@ 2.1 Task Formulation
We denote FIPO as end-to-end text generation. In the training phase, a local optimizer model $\bm{M}_{o-local}$ is supervisedly fine-tuned to generate an optimized prompt $\bm{\hat{x}}^{o}$:
$\bm{\hat{x}}^{o}=\operatorname*{argmax}_{\bm{M}_{o-local}}p(\bm{\hat{x}}^{o}|\bm{x}^{n},[\bm{\hat{y}}^{n}],[\bm{y}^{n}])$ (1)
based on the naive prompt $\bm{x}^{n}$, optional naive response $\bm{\hat{y}}^{n}$, and optional ground truth $\bm{y}^{n}$. In addition, pairwise chosen optimized prompt $\bm{x}^{o+}$ and rejected optimized prompt $\bm{x}^{o-}$ are provided as labels in training. The optional naive response $\bm{\hat{y}}^{n}$ is generated for the naive prompt $\bm{x}^{n}$ using one neural generator model $\bm{M}_{g*}$:
$\bm{\hat{y}}^{n}=\operatorname*{argmax}_{\bm{M}_{g*}}p(\bm{\hat{y}}^{n}|\bm{x}^{n})$ (2)
While in the testing phase, our ultimate target is to obtain a more superior optimized testing response $\bm{\hat{y}}^{o}_{t}$ than the naive testing response $\bm{\hat{y}}^{n}_{t}$, when applying any testing generator $\bm{M}_{g}$ to the optimized testing prompt $\bm{\hat{x}}^{o}_{t}$ and the naive testing prompt $\bm{x}^{n}_{t}$, respectively:
$\bm{\hat{y}}^{o}_{t}\succ\bm{\hat{y}}^{n}_{t}$ (3)
$\bm{\hat{y}}^{o}_{t}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{o}_{t}|\bm{\hat{x}}^{o}_{t})\mbox{, }\bm{\hat{y}}^{n}_{t}=\operatorname*{argmax}_{\bm{M_{g}}}p(\bm{\hat{y}}^{n}_{t}|\bm{\hat{x}}^{n}_{t})$ (4)
where $\bm{M}_{g}$ could be either same as or different from $\bm{M}_{g*}$. And specifically, $\bm{\hat{x}}^{o}_{t}$ is enhanced from $\bm{x}^{n}_{t}$ by the fine-tuned optimizer $\bm{M}_{o-local}$:
$\bm{\hat{x}^{o}_{t}}=\operatorname*{argmax}_{\bm{M_{o}}}p(\bm{\hat{x}}^{o}_{t}|\bm{x}^{n}_{t})$ (5)
In contrast, former ad-hoc APO has no training phase but only the iterative online testing pipeline with mandatory in-box testing response $\bm{\hat{y}}^{o_{i}}_{t}$:
$\bm{\hat{x}}^{o_{i+1}}_{t}=\operatorname*{argmax}_{\bm{M}_{o-api}}p(\bm{\hat{x}}^{o_{i+1}}_{t}|\bm{x}^{o_{i}}_{t},\bm{\hat{y}}^{o_{i}}_{t})\mbox{, }\bm{x}^{o_{1}}_{t}=\bm{x}^{n}_{t}$ (6)
$\bm{\hat{y}}^{o_{i}}_{t}=\operatorname*{argmax}_{\bm{M}_{g-in}}p(\bm{\hat{y}}^{o_{i}}_{t}|\bm{x}^{o_{i}}_{t})\mbox{, }\bm{\hat{y}}^{o_{1}}_{t}=\bm{\hat{y}}^{n}_{t}$ (7)
where $\bm{M}_{g-in}$ is a prior in-box generator.
2.2 Modular Template
As aforementioned in section 2.1, we first design a modular template that ensures flexibility in content management. Figure 2 illustrates our template, shown in the middle, taking mandatory naive task instruction $\bm{x}^{n}$, optional naive response $\bm{\hat{y}}^{n}$ and optional ground truth response $\bm{y}^{n}$ as inputs. Additional description is then appended for clarity, in which we directly claim the optionality of $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$: “The optional sliver response ··· based on the sliver prompt ··· optional golden response ···”.
We use this modular template for all sections in FIPO, including the dataset collection, fine-tuning the local optimizer $\bm{M}_{o-local}$, and testing various downstream generators $\{\bm{M}_{g}\}$. The key difference between these sections is we accordingly adjust the optional responses, thus addressing potential exposure bias between the training and the inference phases (Eq. 1 vs. Eq. 5): (1) Keep both in data collection. We introduce the collection of our POP data in section 2.3; (2) Diversely keep partial responses in training. We present this strategy in section 2.4; (3) Remove all responses in testing. We report testing results in section 3.2.
Figure 3: Step 3 of FIPO: transitional dataset diversification and several mainstream fine-tuning strategies.
2.3 Prompt Optimization Preference Data
@@
We decide to distill the refined prompt optimization capabilities from prominent yet proprietary LLMs, instead of directly integrating them in an ad-hoc way. Thus, we collect the Prompt Optimization Preference (POP) data. Shown in Figure 2, to ensure the most directional optimization, we send naive prompt $\bm{x}^{n}$, naive response $\bm{\hat{y}}^{n}$, and ground truth response $\bm{y}^{n}$ to one suboptiomal LLM GPT-3.5-turbo and one optimal LLM GPT-4222https://platform.openai.com/docs/models, to collect contrastive POP data ($\bm{x}^{o+}$, $\bm{x}^{o-}$). The naive prompt $\bm{x}^{n}$ is sampled from the Alpaca dataset, which contains 52K diverse instructions and corresponding responses $\bm{\hat{y}}^{n}$ generated by the Text-davinci-003 model Taori et al. (2023). We also collect another GPT4-generated response for the Alpaca dataset from public literature Ghosal et al. (2023). There are no official ground truth responses (e.g., from human experts) for the Alpaca data, we therefore consider GPT-4 responses as the ground truth response $\bm{y}^{n}$, given its demonstrated analytical capabilities comparable to humans Pan et al. (2023a). As shown at the bottom of Figure 2, GPT-4 offers a more pedagogical step-by-step optimized prompt compared to GPT-3.5-turbo. We report complete collection template in Table 10.
@@ 2.4 Fine-tuning Strategies
We introduce our fine-tuning strategies in Figure 3, consisting of an initial step of transitional dataset diversification followed by strategic fine-tuning.
Dataset Diversification. In the left hand of Figure 3, we evenly split the 30k samples as eight types depending on the existence of naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$, as well as a format condition “generation” or “multi-choice”. Directionally fine-tuning optimizer has to relay on pre-generated responses, while no any response will be exposed during inference. Hence, the dataset diversification is necessary to help reduce the exposure gap between training and testing, and generalize the original “generation” instruction format to another common “multi-choice” instruction format. The left bottom corner in Figure 3 takes an example of Type 6. The responses $\bm{\hat{y}}^{n}$ and $\bm{y}^{n}$ are modified as candidates adhere to the naive prompt $\bm{x}^{n}$. We then set “A” and “B” as new naive response $\bm{\hat{y}}^{n}$ and ground truth response $\bm{y}^{n}$ [^3] More details of data diversification are in Appendix A..
@@
On the other hand, the right middle shows a contrastive fine-tuning methodology: Preference Optimization, such as Direct Preference Optimization (DPO) Rafailov et al. (2023) and Identity Preference Optimization (IPO) Azar et al. (2023). Preference Optimization takes pairwise rejected label $\bm{x}^{o-}$ and chosen label $\bm{x}^{o+}$ as supervision. One of the core differences from Preference Optimization and SFT is the former one not only encourage the generation of optimal preference, but also dampen the generation of suboptimal preference:
$L_{DPO}(M_{o})=-\mathbb{E}_{(x^{n},\hat{y}^{n},y^{n},x^{o+},x^{o-})\sim D}[\log\sigma(\beta\cdot\Delta)]$ (9)
$L_{IPO}(M_{o})=-\mathbb{E}_{(x^{n},\hat{y}^{n},y^{n},x^{o+},x^{o-})\sim D}[\Delta-\frac{1}{2\beta}]^{2}$ (10)
$\Delta=\log\frac{M_{o}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o+}|x^{r},\hat{y}^{r},y^{r})}-\log\frac{M_{o}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}{M_{ref}(x^{o-}|x^{r},\hat{y}^{r},y^{r})}$ (11)
where $\bm{\beta}$ is a hyperparameter factor. $\bm{M_{ref}}$ refers to the reference model, which is a frozen copy of initial weights of $\bm{M_{o}}$. The equations indicate that IPO is a regularized version of DPO as it limits the optimization range with squares.
Additionally, inspired by self-updating alignment Lee et al. (2023); Anthropic (2022); Yuan et al. (2024), we develop a Iterative Preference Learning (IPL) strategy for self-rewarding prompt optimization. After each iteration of prompt optimization, we ask the optimizer itself to determine if it successfully generate a superior prompt $\bm{x}^{n+}$ with a better response $\bm{\hat{y}}^{n+}$, and if so, to automatically replace the previous inferior prompt $\bm{x}^{n}$ and response $\bm{\hat{y}}^{n}$, leading to more rigorous training in next iteration:
$L_{IPL}(M_{o})=-\mathbb{E}_{(x^{n+},\hat{y}^{n+},y^{n},x^{o+},x^{o-})\sim D}G(\Delta)$ (12)
$x^{n+}=M_{o}(x^{n})\mbox{, }\hat{y}^{n+}=M_{o}(x^{n+})$ (13)
$x^{n+}=\left\{\begin{array}[]{l}x^{n+}\mbox{, }M_{o}(x^{n+},y^{n})\succ M_{o}(x^{n},y^{n})\\ x^{n}\mbox{, otherwise}\end{array}\right.$ (14)
where $G(*)$ denotes as either IPO or DPO loss [^4] Algorithmic details of IPL lie in Appendix A and  D..